Dear Richard dear all,
please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:
Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini, S. Esposito, C. Arcidiacono, A. Riccardi, M. Xompero, R. Briguglio
ADS: E. Anaclerio
MG: R. Biasi
Total # of person: 12 W unit
1) Final routing of WFS board cabling to accommodate the fabric cover has ben done.
2) WFS optical alignment
a) Pupil geometry has been found out of specs by 0.2 pixel (36.2 instead of 36 pixel measured as side of the pupil center square
b) A possible malfunctioning of the fast modulating tip-tilt mirror is suspected. Further investigation will be done tomorrow.
c) The cover holder mounted on lens L1 has been refurbished to avoid collision with the telecentric lens mounted in front of the WFS stage.
3) CCD47 The entrance window of the CCD47 has been cleaned up to remove some blurriness in the images. However the operation did not solve the problem. We suppose some air trapped in the CCD47 optical window closure has been condensing on the CCD internal window. This require to access the CCD47 window to clean it out. AO control SW
1) The SVN repository @ Tucson has been populated. It will be used to control Mt.Graham development. A full checkout has been done both on wfsdx and on adsecdx.
2) conf & calib directories have been moved to adsecdx.
3) More work, in conjunction with Jim Davis from Tucson, to configure the AdSec
1) The cooling lines of the electronic cabinet has been tested to for leakage.
2) The electronic cabinet has been powered on and basic communication tested positively..
3) The small box containing the thin shell TS3 has been removed from the main box and placed in the optical lab.
Next update tomorrow.
All the best
- 02 Apr 2010